inverse of 1 2 3

4 5 3

7 8 9


1
Expert's answer
2022-05-15T17:12:46-0400

The inverse of matrix A can be computed using the inverse of matrix formula, by dividing the adjoint of a matrix by the determinant of the matrix.

For a matrix A, its inverse

"A^{-1}=1\/|A| Adj A"

"|A|=1(-1)^{1+1}\begin{vmatrix}
 5 & 3 \\
 8 & 9
\end{vmatrix}
+2(-1)^{1+2}\begin{vmatrix}
 4 & 3 \\
 7 & 9
\end{vmatrix}
+3(-1)^{1+3}\begin{vmatrix}
 4 & 5 \\
 7 & 8
\end{vmatrix} ="

"=45-24-2(36-21)+3(32-35)=21-72+42+96-105=-18"

AdjA=Transpose of its cofactor matrix

"C=\begin{pmatrix}
 +\begin{vmatrix}
 5 & 3 \\
 8 & 9
\end{vmatrix}& -\begin{vmatrix}
 4 & 3 \\
 7 & 9
\end{vmatrix} & +\begin{vmatrix}
 4 & 5 \\
 7 & 8
\end{vmatrix} \\
 -\begin{vmatrix}
 2& 3 \\
 8 & 9
\end{vmatrix}& +\begin{vmatrix}
 1 & 3 \\
 7 & 9
\end{vmatrix} & -\begin{vmatrix}
 1 & 2 \\
 7 & 8
\end{vmatrix}\\
 +\begin{vmatrix}
 2& 3 \\
 5 & 3
\end{vmatrix} & -\begin{vmatrix}
 1 & 3 \\
 4 & 3
\end{vmatrix} &+\begin{vmatrix}
 1 & 2 \\
 4 & 5
\end{vmatrix}
\end{pmatrix}" "=\begin{pmatrix}
 21 & -15 &-3 \\
 6 & -12&6\\
-9 & 9&-3
\end{pmatrix}"

"AdjA=\begin{pmatrix}
 21 & 6 &-9\\
 -15 & -12&9\\
 -3&6&-3
\end{pmatrix}"


Thus, "A^{-1}= \begin{pmatrix}
 \frac {21}{-18} & \frac 6 {-18}&\frac {-9} {-18}\\
 \frac {-15} {-18} & \frac {-12} {-18}&\frac 9 {-18}\\
 \frac {-3} {-18}&\frac 6 {-18}& \frac {-3} {-18}
\end{pmatrix}=\begin{pmatrix}
 -\frac 7 6 & -\frac 1 3 & \frac 1 2 \\
 \frac 5 6 & \frac 2 3 & -\frac 1 2 \\
 \frac 1 6 & -\frac 1 3 & \frac 1 6
\end{pmatrix}"